if you have a 10 Malary sodium hydroxide solution, how could you prepare 50 mL of a three Malary sodium hydroxide solution?

malary should say molarity
Answer #1

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 10 M

M2 = 3 M

V2 = 50 mL

use:

M1*V1 = M2*V2

V1 = (M2 * V2) / M1

V1 = (3*50)/10

V1 = 15 mL

Take 15 mL of 10 Molar solution and add enough water until final volume is 50 mL